# Upper and Lower Bounds for the Linear Ordering Principle

### v1

@@ 3 $\mathbf{{P}}^{\mathbf{{prMA}}}\subseteq\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$
In this section we prove Theorem 1 by, essentially, expanding the proof of $\mathbf{{MA}}\subseteq\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ in [KP24]. We use the following statements from that paper and an earlier paper by Korten [Kor21].
Definition 3.1 ([Kor21, Definitions 7, 8]).
@@
Proposition 3.2 ([Kor21, Theorem 3]).
PRG reduces in polynomial time to Avoid.
@@
4 Which Karp–Lipton–style Collapse is Better?
Chakaravarthy and Roy proved two Karp–Lipton–style collapses: down to $\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ [CR06] and down to $\mathbf{{P}}^{\mathbf{{prMA}}}$ [CR11]. These two classes seem to be incomparable thereby rising the question: which collapse result is stronger? We observe that the collapse to $\mathbf{{P}}^{\mathbf{{prMA}}}$ can actually be deepened to $\mathbf{{P}}^{\mathbf{{prOMA}}}$, where $\mathbf{{prOMA}}$ is the oblivious version of $\mathbf{{prMA}}$ — and subsequently show that latter class is contained in both previous classes. That is, $\mathbf{{P}}^{\mathbf{{prOMA}}}\subseteq\mathbf{{P}}^{\mathbf{{prMA}}}\cap\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$. Indeed, the “internal collapse” of $\mathbf{{prMA}}$ (and, in fact, even $\mathbf{{prAM}}$) to $\mathbf{{prOMA}}$, under the assumption that $\mathbf{{NP}}\subseteq\mathbf{{P/poly}}$, is implicit in [AKSS95]. Nonetheless, we include a formal proof in Appendix A in order to present a self-contained argument.
Proposition 4.1 (Prop. A.1).
@@ 4 Which Karp–Lipton–style Collapse is Better?
  1. 1.
  2. 2.
The first inclusion is essentially proved in [CR06, Theorem 3], which says that $\mathbf{{MA}}\subseteq\mathbf{{NO_{2}^{\mathbf{{\scriptscriptstyle P}}}}}\ (\subseteq\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}})$. One needs to notice that the proof goes through for promise classes.
Proposition 4.2 (Prop. B.1).
We include a formal proof of this proposition in Appendix B.
The second inclusion seems novel and is proved in the following lemma:
Lemma 4.3.
@@
$W_{n}:=(w^{(0)}_{1},\ldots,w^{(0)}_{p(n)},w^{(1)}_{1},\ldots,w^{(1)}_{p(n)})$
$a:=A(q,u^{(1)}_{|q|},v^{(0)}_{|q|}),\qquad$ $b:=A(q,v^{(1)}_{|q|},u^{(0)}_{|q|}),$
$c:=A(q,u^{(1)}_{|q|},u^{(0)}_{|q|}),\qquad$ $d:=A(q,v^{(1)}_{|q|},v^{(0)}_{|q|}),$
  • Suppose $x\in L$ and $U=W_{\left|{x}\right|}$.
  • Suppose $x\not\in L$ and $V=W_{\left|{x}\right|}$.
Theorem 4.4 (Theorem 3, restated).
Proof.
By Proposition 4.2 and Lemma 4.3. ∎
Remark 4.5.
In this section we prove Theorem 4. Our proof strategy is as follows: Given a point in a linear order, we aim to move “down the order” (i.e. towards “smaller” points). At each stage we will skip over a constant fraction of the points remaining on our way to the minimum. In order to find the next point, we will employ a binary-search-like procedure to determine the bits of the desired point, one coordinate at a time. Here is where our $\mathbf{{prSBP}}$ oracle comes into play: At each step, we look at the remaining set of points partitioned into two subsets: the points where the appropriate bit is 0 and where that bit is 1, and select the subset with the (approximately) smaller average rank.
5.1 Approximate Counting
For a unary relation $R(x)$, we denote the number of elements in $R$ as #$\#_{x}R(x):=\left|{{\left\{x\;\left|\;x\in R\right.\right\}}}\right|$. We will also abbreviate this notation to #$\#_{x}R$. For $k\in\mathbb{N}$, we define $R^{\otimes k}$ — the $k$-th tensor power of $R$ as
Observation 5.1.
Then #$\#_{x}(R^{\otimes k})=(\#_{x}R)^{k}$.
Note that queries to $\mathbf{{prSBP}}$ can be thought of as queries to a specific promise problem.
Definition 5.2 (Set-Size Estimation, SSE and WSSE).
Lemma 5.3 (Implicit in [BGM06]).
SSE is $\mathbf{{prSBP}}$-complete.
Lemma 5.4 (Implicit in [GS86]).
WSSE is $\mathbf{{prAM}}$-complete.
In particular, these complete problems showcase that $\mathbf{{prSBP}}\subseteq\mathbf{{prAM}}$.
Observation 5.5 (Useful inequalities).
Let $x,y\geq 0$ and $0\leq\varepsilon\leq 1$. Then
  1. 1.
    $4^{\varepsilon/3}\leq 1+\varepsilon$ (Bernoulli’s inequality),
  2. 2.
    $\frac{x+\varepsilon y}{x+y}\geq\varepsilon$.
Next, we show how to approximate deterministically the number of satisfying assignments for a Boolean circuit using a $\mathbf{{prSBP}}$ oracle (i.e. $\mathbf{{FP}}^{\mathbf{{prSBP}}}$). Previously, it was shown that this task could be carried out by a randomized algorithm using an $\mathbf{{NP}}$ oracle ($\mathbf{{FBPP}}^{\mathbf{{NP}}}$) [Sto85, JVV86] and by a deterministic algorithm using a $\mathbf{{prAM}}$ oracle ($\mathbf{{FP}}^{\mathbf{{prAM}}}$)  [^10] Essentially using an oracle access to WSSE. [Sip83, GS86].
Lemma 5.6.
#$\#_{x}C\leq t\leq 4^{\varepsilon/3}\cdot\#_{x}C\leq(1+\varepsilon)\#_{x}C$
in time polynomial in $n$, the size of $C$ and $\frac{1}{\varepsilon}$, given oracle access to SSE.
Proof.
  1. 1.
    If $\left(C,1\right)\not\in O$  return $t=0$triangleright #$\#_{x}C=0$
  2. 2.
  3. 3.
    return $t=\left\lfloor{2^{i/k}}\right\rfloor$
If #$\#_{x}C=0$ then $\left(C,1\right)\not\in O$ and the algorithm outputs $t=0$. Otherwise, #$\#_{x}\hat{C}\geq 1$ and hence $\left(\hat{C},1\right)\in O$. On the other hand, by definition #$\#_{x}\hat{C}\leq 2^{nk}$ and hence $\left(\hat{C},2^{nk+1}\right)\not\in O$. Therefore, $i$ is well-defined and we have that:
In conclusion,
#$\#_{x}C\leq t\leq 4^{1/k}\cdot\#_{x}C\leq 4^{\varepsilon/3}\cdot\#_{x}C\leq(1+\varepsilon)\#_{x}C.$
5.2 Estimating the Average Rank w.r.t. a Linear Order
[…]
Observation 5.7.
[…]
Fix any strict linear order $<$ on $U$.
Definition 5.8.
For an element $\alpha\in U$ we define its rank as $\mathrm{rank}(\alpha):=\left|{{\left\{x\in U\;\left|\;x<\alpha\right.\right\}}}\right|$. We can extend this definition to non-empty subsets $S\subseteq U$ of $U$ by taking the average rank: define $\mathrm{rank}(S):=\frac{\sum_{x\in S}\mathrm{rank}(x)}{\left|{S}\right|}$. If $S={\left\{x\in U\;\left|\;C(x)=1\right.\right\}}$ is described by a circuit $C$, we use the same notation: $\mathrm{rank}(C)=\mathrm{rank}(S)$.
Below are some useful observations that we will use later.
Observation 5.9.
  • For a non-empty subset $S\subseteq U$, |$\left|{{\left\{(\upsilon,\alpha)\in U\times S\;\left|\;\upsilon<\alpha\right.\right\}}}\right|=\left|{S}\right|\cdot\mathrm{rank}(S).$
  • Let $S_{0},S_{1}\subseteq U$ be two non-empty disjoint subsets of $U$. Then
[…]
Lemma 5.10.
Proof.
[…]
#$\#_{x}C\leq t_{C}\leq 4^{\varepsilon}\cdot\#_{x}C\text{ and }\#_{(x,y)}D\leq t_{D}\leq 4^{\varepsilon}\cdot\#_{(x,y)}D.$
Therefore we obtain:
5.3 Finding the Minimum Using a $\mathbf{{prSBP}}$ Oracle
We use the approximation algorithms developed above in order to find an element that is much closer to the minimum than a given element. The following lemma describes the procedure Back that given an element $\alpha$ finds another element $\beta$ whose rank is smaller by a constant factor. We will use this procedure afterwards in order to find the minimum in a polynomial number of iterations.
[…]
Lemma 5.11.
Proof.
Consider the following procedure:
$\textsc{Back}(E,\alpha):$
  1. 1.
[…]
  2. 2.
    Set $\varepsilon=1/(8n)$.
  3. 3.
    For $i=1$ to $n$:
    1. (a)
      For $b\in{\left\{0,1\right\}}$: define $C_{b}:=C|_{x_{i}=b}$
    2. (b)
      For $b\in{\left\{0,1\right\}}$: if #$\#_{x}C_{b}=0$ then set $C:=C_{1-b},\ \beta_{i}:=1-b$; continue to the next $i$
      $\triangleright$ If one of the sets is empty, we choose the other one
    3. (c)
    4. (d)
      If $r_{1}\geq r_{0}$ then $C:=C_{0},\beta_{i}:=0$ else $C:=C_{1},\beta_{i}:=1$
      $\triangleright$ Choose the set with smaller approximate order
After each iteration one more variable $x_{i}$ gets its value $\beta_{i}$ and is substituted into $C$, that is, in the current circuit $C$ variables $x_{1},\ldots,x_{i}$ are replaced by the corresponding constants $\beta_{1},\ldots,\beta_{i}$. We claim that after each iteration the rank of the resulting circuit is bounded from the above: $\mathrm{rank}(C)\leq 4^{2\varepsilon i}\cdot\frac{\mathrm{rank}(\alpha)}{2}.$
Indeed, by Observation 5.9, before the first iteration, we have that $\mathrm{rank}(C)=\frac{\mathrm{rank}(\alpha)}{2}$. Now consider any iteration. If $C_{1}$ or $C_{0}$ are empty, then $\mathrm{rank}(C)$ remains the same and $4^{2\varepsilon i}\leq 4^{2\varepsilon(i+1)}$. Otherwise, by Lemma 5.10, for $b\in{\left\{0,1\right\}}:$
If $r_{1}\geq r_{0}$ then $\mathrm{rank}(C_{1})\geq r_{1}\cdot 4^{-\varepsilon}\geq r_{0}\cdot 4^{-\varepsilon}\geq\mathrm{rank}(C_{0})\cdot 4^{-2\varepsilon}$ and therefore by Observations 5.9 and 5.5(2)
Equivalently, $\mathrm{rank}(C_{0})\leq\mathrm{rank}(C)\cdot 4^{2\varepsilon}$. Similarly, if $r_{1}<r_{0}$ then $\mathrm{rank}(C_{1})\leq\mathrm{rank}(C)\cdot 4^{2\varepsilon}.$ Therefore, at each step the rank is multiplied at most by $4^{2\varepsilon}$.
Observation 5.12.
Proof.
We are now ready to prove the main result of this section.
Theorem 5.13.
Proof.
  1. 1.
    Check that $E$ is indeed a linear order using Observation 5.7.
  2. 2.
    Let $\alpha:=0^{n},\ \beta:=1^{n}$.
  3. 3.
  4. 4.
    Output $\alpha$.
Given Observation 5.7, we can assume w.l.o.g. that $E$ is a linear order circuit. We claim that the algorithm will output the minimal element after at most $2n$ iterations. Indeed, by Lemma 5.11, the rank of the element $\alpha$ after $2n$ iterations satisfies $\mathrm{rank}(\alpha)\leq\frac{\mathrm{rank}(1^{n})}{\sqrt{2}^{2n}}\leq\frac{2^{n}-1}{2^{n}}<1$, thus the “While” cycle will be terminated before that. ∎
6 Further Research
@@
    Chakaravarthy and Roy [CR11] asked whether $\mathbf{{P}}^{\mathbf{{prMA}}}$ and $\mathbf{{P}}^{\mathbf{{prS_{2}^{\mathbf{{\scriptscriptstyle P}}}}}}$ are contained in $\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$. While we resolved the first question, the second one remains open. We note that, while in the input-oblivious world both inclusions hold ($\mathbf{{P}}^{\mathbf{{prOMA}}}\subseteq\mathbf{{P}}^{\mathbf{{prO_{2}^{\mathbf{{\scriptscriptstyle P}}}}}}\subseteq\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$, Theorem 4.4), the proof of the latter inclusion (Lemma 4.3) is essentially input-oblivious (one needs to give all the certificates for the oracle non-adaptively, and queries cannot be predicted because oracle answers cannot be predicted for promise problems).
@@
    While Proposition 4.1 shows a collapse of $\mathbf{{PH}}$ to $\mathbf{{P}}^{\mathbf{{prOMA}}}$, it does not immediately imply fixed-polynomial size lower bounds for $\mathbf{{P}}^{\mathbf{{prOMA}}}$, since this class is unlikely to contain $\mathbf{{NP}}$, so in the case that $\mathbf{{NP}}\not\subseteq\mathbf{{P/poly}}$ it may indeed miss hard languages. One can observe that Santhanam’s proof [San09] of $\mathsf{Size}[n^{k}]$ lower bounds for promise problems in $\mathbf{{prMA}}$ is input-oblivious. Indeed, the presented hard promise problems are actually in $\mathbf{{prOMA}}$! However, it does not yield a language in $\mathbf{{P}}^{\mathbf{{prOMA}}}$ that is hard for $\mathsf{Size}[n^{k}]$, and we leave this question for further research.
@@
    Similarly to $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$, one could define a class of languages reducible to Avoid. A similar class of search problems, $\mathbf{{APEPP}}$, has been defined by [KKMP21, Kor21] (and Korten [Kor21] proved that constructing a hard truth table is a problem that is complete for this class under $\mathbf{{P}}^{\mathbf{{NP}}}$-reductions); however, we are asking about a class of languages. Korten and Pitassi have shown that $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ can be equivalently defined using many-one, Turing, or $\mathbf{{P}}^{\mathbf{{NP}}}$-reductions, thus there are several options. One can observe that the containment $\mathbf{{P}}^{\mathbf{{prMA}}}\subseteq\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ (Theorem 3.4) is essentially proved via the intermediate class $\mathbf{{P}}^{\textsc{Avoid},\mathbf{{NP}}}$ that uses both an oracle for Range Avoidance (a single-valued or an essentially unique [KP24] version) and an oracle for SAT. Can one prove that one of the containments in $\mathbf{{P}}^{\mathbf{{prMA}}}\subseteq\mathbf{{P}}^{\textsc{Avoid},\mathbf{{NP}}}\subseteq\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ is in fact an equality?
@@
    As was mentioned, the $\mathbf{{FP}}^{\mathbf{{prSBP}}}$ procedure for approximate counting can be implemented using parallel (i.e. non-adaptive) oracle queries. On the other hand the containment $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}\subseteq\mathbf{{P}}^{\mathbf{{prSBP}}}$, which uses approximate counting as a black-box subroutine, seems to require sequential, adaptive queries. Could one implement the latter containment using parallel queries? In particular, as $\mathbf{{PP}}$ is consistent with $\mathbf{{prSBP}}$ [BGM06] and is closed under non-adaptive Turing reductions [FR96], this would imply that $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}\subseteq\mathbf{{PP}}$. Note that it is unknown even whether $\mathbf{{P}}^{\mathbf{{NP}}}\subseteq\mathbf{{PP}}$, moreover, there is an oracle separating these two classes [Bei94].

### v2

@@ 5 $\mathbf{{P}}^{\mathbf{{prMA}}}\subseteq\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$
In this section we prove Theorem 1 by, essentially, expanding the proof of $\mathbf{{MA}}\subseteq\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ in [KP24]. We use the following statements from that paper and an earlier paper by Korten [Kor22]. We then proceed to the input-oblivious setting.
5.1 The non-input-oblivious setting
Definition 5.1 ([Kor22, Definitions 6, 7], [Kor21, Definitions 7, 8]).
@@ 5.1 The non-input-oblivious setting
Proposition 5.2 ([Kor22, Theorem 2], [Kor21, Theorem 3]).
PRG reduces in polynomial time to a single Avoid query.
@@
5.2 The Input-Oblivious Setting
Korten proves (Prop. 5.2) that PRG (Def. 5.1) reduces to Avoid, and Korten and Pitassi [KP24] compute Avoid in $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}=\mathbf{{P}}^{\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}}$. Since PRG has a unary input, one can observe that PRG can be computed using an input-oblivious oracle. This gives raise to tighter containments. In the non-input oblivious setting they are equalities, but input-oblivious classes lack some of the nice closureness properties and need a separate treatment.
Proposition 5.5.
PRG can be computed in deterministic polynomial time with an $\mathbf{{OL_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ oracle.
Proof.
It is shown in [Kor22] (see Prop. 5.2) that PRG consisting of $m=n^{6}$ strings can be computed in deterministic polynomial time with a single oracle query to Avoid; denote the algorithm generating this query by $K$, let us assume w.l.o.g. that it outputs a circuit of bit size $n^{d}$ with $n^{c}$ inputs, for certain integer constants $d>c>2$. Since the input to this problem is unary, for each $n$, a single instance $C_{n}$ of Avoid is to be solved, and its solution (i.e. any string outside of the image of $C_{n}$) solves PRG (in fact, without further processing). A specific solution to Avoid can be computed using a deterministic polynomial-time truth-table reduction to a language $L\in\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ [KP24], that is, there is a polynomial $p$, a $p$-time DTM $R$ computing the reduction, a $p$-time DTM $T$ computing the truth table, a language $L\in\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ with a polynomial-time verifier $V$ computing an order $_{V}$ such that $T[a_{1},a_{2},\ldots a_{r}]$ is a correct solution to Avoid, where $r$ is the number of queries made by $R$ (w.l.o.g. they are of the same size), and for each $i$, $a_{i}$ is the first bit of the minimum certificate wrt $_{V}$ applied to the $i$-th query.
Our oracle language $L^{\prime}\in\mathbf{{OL_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ is verified by the following verifier $U$ (defining its order $_{U}$), which merges queries using a construction similar to [KP24].
For two bit strings $\alpha,\beta$ we define the relation: $\alpha\mathrel{<_{lex}}\beta$ to be 1 iff $\alpha$ is lexicographically smaller than $\beta$ (and 0, otherwise).
Input: $w$
Certificates: $y$, $z$
Algorithm:
  1. 1.
    If $y=z$ as bit strings, then return 0. // $y\mathrel{\nless_{U}}z$
  2. 2.
    Compute $t:=|w|$.
  3. 3.
    Compute $n:=\lfloor\sqrt[d]{|w|}\rfloor$.
  4. 4.
    Compute i:=twi:=t-w.
  5. 5.
    Let $a$ be the first bit of $y$, and $b$ be the first bit of $z$.
    // Informally, this bit is a claim for the value of the $i$-th bit of PRG.
  6. 6.
    If $i>n^{c}$ then // Out of range.
    WWWif $a\neq b$, then return the result of comparison $a<b$, otherwise return $y\mathrel{<_{lex}}z$.
  7. 7.
    Run $K(1^{n})$, denote the resulting circuit by $C_{n}$.
    // Note that the length of its description is $n^{d}$.
  8. 8.
    Run $R(C_{n})$ to compute queries $q_{1},q_{2},\ldots,q_{r}$.
  9. 9.
    Parse $y$ as $axy_{1}\ldots y_{r}$ and $z$ as $bx^{\prime}z_{1}\ldots z_{r}$, where $a,b\in\{0,1\}$, $x,x^{\prime}\in\{0,1\}^{n^{d}}$$y_{1},\ldots,y_{r},z_{1},\ldots,z_{r}\in\{0,1\}^{s}$, where $s$ is the bit size of elements of $_{V}$ for the queries computed by $R$.
    // $x,x^{\prime}$ are candidates for the solution of Avoid computed by $R,T$.
  10. 10.
    Call $y$ syntactically incorrect if either $x[i]$ (the $i$-th bit of $x$) differs from $a$ or $T[y_{1}[1],\ldots,y_{r}[1]]\neq x$.
    // That is, $y$ does not claim that the answer is $x[i]$, or its sub-certificates do not yield $x$ as $T$’s answer.
    Define syntactically incorrect $z$ similarly.
  11. 11.
    If both $y$ and $z$ are syntactically incorrect, return $y\mathrel{<_{lex}}z$.
  12. 12.
    If exactly only one of $y,z$ is syntactically incorrect, state that it is greater than the other certificate, return $1$ or 0 accordingly.
  13. 13.
    Now $y,z$ are syntactically correct.
    If $y_{1}\ldots y_{r}\neq z_{1}\ldots z_{r}$ then find the first $j$ such that $y_{j}\neq z_{j}$ and return $V(q_{j},y_{j},z_{j})$.
  14. 14.
    We are done and never get to this point: the certificates are different, but the sub-certificates are equal, thus $x\neq x^{\prime}$ so one of the certificates had to be recognized as syntactically incorrect by $T$ lookup.
Our verifier $U$ is input-oblivious (it simply ignores the input and only uses its length), it computes a linear order, and its minimal element starts with the bit equal to the $i$-th bit of the pseudorandom generator computed by the composition of $K$, $R$, and $T$. Thus $U$ defines a language in $\mathbf{{OL_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$, and consequent queries $1^{n^{d}+i}$ to it reveal the bits of the pseudorandom generator. ∎
Corollary 5.6.
  1. 1.
    $\mathbf{{P}}^{\mathbf{{prMA}}}\subseteq\mathbf{{P}}^{\mathbf{{OL_{2}^{\mathbf{{\scriptscriptstyle P}}}}},\mathbf{{NP}}}\subseteq\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$.
  2. 2.
    $\mathbf{{P}}^{\mathbf{{prOMA}}}\subseteq\mathbf{{P}}^{\mathbf{{OL_{2}^{\mathbf{{\scriptscriptstyle P}}}}},\mathbf{{prONP}}}\subseteq\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}\cap\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$.
Proof.
The first inclusion follows similarly to the proof of Theorem 5.4: the lower DTM can compute PRG using its $\mathbf{{OL_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ oracle (Prop. 5.5) and then use it to derandomize $\mathbf{{prMA}}$ (resp., $\mathbf{{prOMA}}$) using its $\mathbf{{NP}}$ (resp., $\mathbf{{prONP}}$) oracle.
@@ 5.2 The Input-Oblivious Setting
5.3 An even better Karp–Liption–style collapse?
One can define a promise version of $\mathbf{{prOMA}}$ where the machine satisfies the bounded-error promise on the promise set but it has to satisfy the input-oblivious promise everywhere.
Definition 5.7.
A promise problem $\Pi=(\Pi_{\textsc{\tiny YES}},\Pi_{\textsc{\tiny NO}})$ belongs to $\mathbf{{OprMA}}$ if there is a polynomial-time deterministic Turing machine $A$ and, for every $n\in\mathbb{N}$, there exists $w_{n}$ (a witness that serves for every positive instance of length $n$), that satisfy the following conditions:
  • If $x\in\Pi_{\textsc{\tiny YES}}\cap\{0,1\}^{n}$, then $\forall r\ A(x,r,w_{n})=1$,
  • If $x\in\Pi_{\textsc{\tiny NO}}$, then $\forall w\ \Pr_{r}[A(x,r,w)=1]<1/2$,
  • If $x\notin\Pi_{\textsc{\tiny YES}}$, then $\forall w\ \Pr_{r}[A(x,r,w)=1]<1$.
We conjecture that by careful inspection of the proof of [CR11] one can check that the only reason to use promise problems is a lack of guarantee for bounded probability of error, therefore:
Conjecture 5.8.
The Karp–Lipton–style collapse of Proposition 4.1 to $\mathbf{{P}}^{\mathbf{{prOMA}}}$ is actually to $\mathbf{{P}}^{\mathbf{{OprMA}}}$.
Anyway, this definition gives rise to an even nicer corollary than Corollary 5.6.
Corollary 5.9.
$\mathbf{{P}}^{\mathbf{{OprMA}}}\subseteq\mathbf{{P}}^{\mathbf{{OL_{2}^{\mathbf{{\scriptscriptstyle P}}}}},\mathbf{{ONP}}}\subseteq\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}\cap\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$.
@@ 5.3 An even better Karp–Liption–style collapse?
The second inclusion is already proved in Corollary 5.6, we need to show only the first one. It’s also analogous: similarly to the proof of Theorem 5.4: the lower DTM can compute PRG using its $\mathbf{{OL_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ oracle (Prop. 5.5) and then use it to derandomize $\mathbf{{OprMA}}$ using its $\mathbf{{ONP}}$ oracle (we do not need a $\mathbf{{prONP}}$ oracle here, because the promise in $\mathbf{{OprMA}}$ does not concern the input-obliviousness). ∎
6 Discussion and Further Research
Some Unresolved Containments
@@ 6 Discussion and Further Research
    Chakaravarthy and Roy [CR11] asked whether $\mathbf{{P}}^{\mathbf{{prMA}}}$ and $\mathbf{{P}}^{\mathbf{{prS_{2}^{\mathbf{{\scriptscriptstyle P}}}}}}$ are contained in $\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$. While we resolved the first question, the second one remains open. We note that, although in the input-oblivious world both inclusions hold ($\mathbf{{P}}^{\mathbf{{prOMA}}}\subseteq\mathbf{{P}}^{\mathbf{{prO_{2}^{\mathbf{{\scriptscriptstyle P}}}}}}\subseteq\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$, Corollary 4.4), the proof of the latter inclusion (Theorem 5) is essentially input-oblivious (one needs to give all the certificates for the oracle non-adaptively, and queries cannot be predicted because oracle answers cannot be predicted for promise problems).
@@
    As was mentioned, the $\mathbf{{FP}}^{\mathbf{{prSBP}}}$ procedure for approximate counting can be implemented in $\mathbf{{FP}}^{\mathbf{{prSBP}}}_{\parallel}$ — that is, using parallel (i.e. non-adaptive) oracle queries. On the other hand the containment $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}\subseteq\mathbf{{P}}^{\mathbf{{prSBP}}}$, which uses approximate counting as a black-box subroutine, seems to require sequential, adaptive queries. Could one implement the latter containment using parallel queries (i.e. show that $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}\subseteq\mathbf{{P}}^{\mathbf{{prSBP}}}_{\parallel}$)? In particular, as $\mathbf{{PP}}$ is consistent with $\mathbf{{prSBP}}$ [BGM06] and is closed under non-adaptive Turing reductions [FR96], this would imply that $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}\subseteq\mathbf{{PP}}$. Note that it is unknown even whether $\mathbf{{P}}^{\mathbf{{NP}}}\subseteq\mathbf{{PP}}$, while $\mathbf{{P}}^{\mathbf{{NP}}}\subseteq\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$. Moreover, there is an oracle separating the former two classes [Bei94].
@@
[…]
@@
    Similarly to $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$, one could define a class of languages reducible to Avoid. A similar class of search problems, $\mathbf{{APEPP}}$, has been defined by [KKMP21, Kor22] (and Korten [Kor22] proved that constructing a hard truth table is a problem that is complete for this class under $\mathbf{{P}}^{\mathbf{{NP}}}$-reductions); however, we are asking about a class of languages. Korten and Pitassi have shown that $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ can be equivalently defined using many-one, Turing, or $\mathbf{{P}}^{\mathbf{{NP}}}$-reductions, thus there are several options. One can observe that the containment $\mathbf{{P}}^{\mathbf{{prMA}}}\subseteq\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ (Theorem 5.4) is essentially proved via the intermediate class $\mathbf{{P}}^{\textsc{Avoid},\mathbf{{NP}}}$ that uses both an oracle for Range Avoidance (a single-valued or an essentially unique [KP24] version) and an oracle for SAT. Can one prove that one of the containments in $\mathbf{{P}}^{\mathbf{{prMA}}}\subseteq\mathbf{{P}}^{\textsc{Avoid},\mathbf{{NP}}}\subseteq\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ is in fact an equality?